the functions f and g are defined as follows
f(x)=4x+2
g(x)=4x^2-x
find f(-3) and g(-4)
simplify your answer as much as possible

Answer by jim_thompson5910(35256)

Plug x = -3 into f(x) to get


f(x)=4x+2


f(-3)=4(-3)+2


f(-3)=-12+2


f(-3)=-10


---------------------


Plug x = -4 into g(x) to get


g(x)=4x^2-x


g(-4)=4(-4)^2-(-4)


g(-4)=4(16)-(-4)


g(-4)=64+4


g(-4)=68